An arrow is shot up with a velocity of 50m/s and the height in meters after t seconds is h=50t-0.36t^2. Find the average velocity over interval [1,1.03]. correct to sour decimal places.
Answer by Alan3354(21608)

An arrow is shot up with a velocity of 50m/s and the height in meters after t seconds is h=50t-0.36t^2. Find the average velocity over interval [1,1.03]. correct to sour decimal places.
-------------
Find the height at 1 second and at 1.03 seconds
h(1) = 50 - 0.36 = 49.64 meters
h(1.03) = 51.5 - 0.381924 = 51.118076 meters
Change in height = 1.478076
--------------
The average speed over the time interval = d/t
= 1.478076/0.03
= 49.2692 m/sec
---------------
I don't know where that formula came from. The 0.36 is a lot less than Earth's gravity.